What is grapichal method of 2x-3y=7; 3x+y=5

1
Expert's answer
2022-03-19T02:40:09-0400

Solution


The two equations "2x-3y=7" and "3x+y=5" are plotted as shown below.





We can see that the two lines intersect at the point "(2, -1)".


Hence the two linear equations "2x-3y=7" and "3x+y=5" give solution "(x,y)=(2,-1)"


Therefore, the graphical method is a method that can be used to find the point of intersection between two non-parallel lines.